OpenStudy (anonymous):

The function f(x) = 6x^2 + 1 is defined over the interval [–2, 5]. If the interval is divided into n equal parts, what is the width of each interval, Δx? 5/n 6/n 7/n 8/n

OpenStudy (anonymous):

@alexray19

OpenStudy (anonymous):

hey!

OpenStudy (michele_laino):

hello!

OpenStudy (anonymous):

So start by figuring out how wide the interval is...

OpenStudy (anonymous):

As in the function's domain, [-2, 5]. How wide is that?

OpenStudy (michele_laino):

for example if I divide your interval by 2, I get:  since the middle point is: (5-2)/2=3/2 so each subset is: 5-3/2=3/2-(-2)=7/2 wide

OpenStudy (anonymous):

a length of 7

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so its c?

OpenStudy (anonymous):

It is C

OpenStudy (anonymous):

give them a medal

OpenStudy (michele_laino):

yes! that's right!